OpenStudy (anonymous):

Assume the given point is on the terminal side of an angle in standard position. Determine the exact values of the six trigonometric functions of the angle: (-2, 10).

OpenStudy (phi):

first, plot the point and draw a line from (0,0) to the point. can you do that ?

OpenStudy (anonymous):

@phi Ok I did it.

OpenStudy (phi):

next, we form a triangle by dropping a line straight down until it hits the x-axis label the sides: x is -2, y is +10 you need the length of the hypotenuse. can you find it ?

OpenStudy (anonymous):

I think so. The length of the legs are 10 and 2 The hypotenuse squared equals the sum of both legs squared. So... 10^2+2^2=104 sqrt104=10.2 So the hypotenuse equals 10.2 Right?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes, but if you want exact numbers, use \( \sqrt{104} = 2 \sqrt{26} \) now form ratios for sin, cos, tan, and their reciprocals c s c , sec , cot

OpenStudy (phi):

notice the question asks for *the exact values*, so we can't use the rounded decimal value for the sqrt(104)

OpenStudy (phi):

use SOH CAH TOA or oscar had a heap of apples to remember the ratios for sin, cos, and tan

OpenStudy (anonymous):

How do I form ratios? :/

OpenStudy (phi):

By definition, the angle is  to find the sin, cos, etc, we use the *reference angle*